# NeuroSim V1.5: Improved Software Backbone for Benchmarking Compute-in-Memory Accelerators with Device and Circuit-level Non-idealities

### v1

@@ Abstract
The exponential growth of artificial intelligence (AI) applications has exposed the inefficiency of conventional von Neumann architectures, where frequent data transfers between compute units and memory create significant energy and latency bottlenecks. Analog Computing-in-Memory (ACIM) addresses this challenge by performing multiply-accumulate (MAC) operations directly in the memory arrays, substantially reducing data movement. However, designing robust ACIM accelerators requires accurate modeling of device- and circuit-level non-idealities. In this work, we present NeuroSim V1.5, introducing several key advances: (1) seamless integration with TensorRT’s post-training quantization flow enabling support for more neural networks including transformers, (2) a flexible noise injection methodology built on pre-characterized statistical models, making it straightforward to incorporate data from SPICE simulations or silicon measurements, (3) expanded device support including emerging non-volatile capacitive memories, and (4) up to 6.5× faster runtime than NeuroSim V1.4 through optimized behavioral simulation. The combination of these capabilities uniquely enables systematic design space exploration across both accuracy and hardware efficiency metrics. Through multiple case studies, we demonstrate optimization of critical design parameters while maintaining network accuracy. By bridging high-fidelity noise modeling with efficient simulation, NeuroSim V1.5 advances the design and validation of next-generation ACIM accelerators. All NeuroSim versions are available open-source at https://github.com/neurosim/NeuroSim.
Index Terms:
Compute-in-memory, AI accelerator, hardware/software co-design, open-source, benchmark
@@ I Introduction
The exponential growth in AI applications has exposed a critical challenge in energy-efficient computing. The fundamental limitation lies in the mismatch between the memory-centric workloads such as in machine learning algorithms and the processing-centric architecture in contemporary hardware platforms. This mismatch causes massive data movement between memory and processing units to consume more than five times the energy of computation [1].
Figure 1: Overview of NeuroSim V1.5: This update revamped the software backbone to support arbitrary neural network models and provide a more intuitive noise simulation workflow tailored to device and circuit experts. NeuroSim V1.5 enables users to evaluate inference accuracy across diverse architectures, including custom convolutional neural networks and transformers, while simulating various device types and noise sources. The platform uniquely supports power, performance, and area (PPA) estimations, facilitating comprehensive system-level design space exploration. Key new features include optimized noise modeling, compatibility with advanced non-volatile memory technologies, and improved runtime performance.
The challenges stem from the data-intensive nature of deep neural networks (DNNs), where a relatively small number of operations is performed per fetched data; this causes system performance to be limited by memory bandwidth, as noted by the roofline model [2]. Compute-in-memory (CIM) has emerged as a promising paradigm to tackle this bottleneck by co-locating data storage and multiply-accumulate (MAC) operations in the same physical arrays, thereby reducing repeated weight fetches. In digital computing-in-memory (DCIM), memory arrays (typically made of modified SRAM bit cells) are augmented with digital multipliers and adder trees for partial-sum accumulation, delivering robust performance with the overhead of adder trees [3].
By contrast, ACIM leverages the intrinsic electrical behavior of memory cells to carry out multiplications directly in the current or charge domain. Specifically, weight values are encoded as conductance (or capacitance), and when an input voltage is applied to a row, the resultant column current (or charge) naturally represents the multiplication of input and weight. Summing all such column outputs yields the MAC result in the analog domain and then being converted back to the digital domain through analog-to-digital converters (ADCs) [4, 5]. A variety of emerging non-volatile memory (NVM) devices have been utilized for these ACIM arrays, including resistive random-access memory (RRAM) [6, 7, 8, 9], phase-change memory (PCM) [10, 11, 12], ferroelectric field-effect transistors (FeFET) [13, 14, 15], and more recently non-volatile capacitors (nvCap) [16, 17, 18]. ACIM implementations can also rely on modified SRAM bit cells, tapping into current-based summation [19, 20, 21, 22, 23, 24], or incorporate on-chip capacitors to accumulate charge [25, 26] or using them for time-domain computing [27]. Several of these memory cells are depicted in a generic CIM macro with peripheral circuits as shown in Fig. 1.
Over the past few years, both DCIM and ACIM have been demonstrated in silicon with memory capacities up to a few megabytes [28]—sufficient for many on-chip inference scenarios such as computer vision tasks involving millions of parameters. DCIM generally offers the robustness of digital processing, making it attractive for high-performance computing, whereas ACIM excels in energy efficiency, an essential requirement for power-constrained edge devices.
ACIM accelerators can theoretically surpass 100 TOPS/W [3, 9, 19, 26]—far exceeding the 1–10 TOPS/W range of contemporary von Neumann platforms—largely because they eliminate most weight-reloading costs by performing the MAC operations directly within the memory arrays. In this paper, one operation is defined as an 8b×8b MAC. The key challenge in ACIM is to manage inherently noisy analog computations to maintain the inference accuracy.
These noise sources span multiple abstraction layers. At the device level, fabrication imperfections lead to device-to-device (D2D) variations and stuck-at-faults. Memory states can also drift over time, degrading compute accuracy [29, 30, 31, 32]. At the circuit level, thermal noise and process variations in peripheral circuits degrade sensing margins, resulting in errors in the ADC processes [33, 18]. At the system level, partial sums may be quantized at a low bit precision, compounding errors as layer outputs are passed from one stage to another [34, 35, 36]. Together, these effects can reduce the inference accuracy of DNNs run on ACIM chips. Consequently, modeling and ultimately mitigating these non-idealities is critical for the design of robust ACIM architectures.
These multi-layered noise effects create complex interactions between device characteristics, circuit design choices, and neural network requirements. Therefore optimizing ACIM accelerators requires systematic exploration across this design space to balance inference accuracy with hardware efficiency. Such an exploration requires simulation frameworks that can evaluate both accuracy and hardware metrics while maintaining practical runtimes even for modern large-scale networks.
Direct SPICE simulation of analog computing arrays is prohibitively expensive for large neural networks, making it impractical to evaluate broad parameter spaces with modern datasets. While tools like CrossSim [37] and AIHWKit [38] offer detailed device modeling and noise-aware training respectively, they lack integrated hardware analysis capabilities and require extensive engineering effort to support new devices or circuits. Other frameworks like CIMLoop [39] focus on hardware metrics without considering accuracy under device and circuit non-idealities. These limitations make it difficult to systematically explore and optimize ACIM designs across both software accuracy and hardware efficiency metrics.
NeuroSim [40] addresses these challenges through an integrated framework that combines behavioral simulation of CIM operation with automated hardware analysis. Building on this foundation, we present NeuroSim V1.5, which introduces several key advances to enable comprehensive design space exploration:
@@
    Seamless integration with TensorRT to support arbitrary pre-trained models, while automating quantization.
@@
    A flexible noise modeling methodology built on pre-characterized statistical models, enabling rapid evaluation of diverse device and circuit designs.
@@
    Support for emerging technologies like charge-based non-volatile capacitive memories (nvCap) and emerging algorithms like vision transformer (ViT).
@@
    Significantly improved runtime performance - up to 6.5× faster than previous versions of NeuroSim.
These capabilities maintain seamless compatibility with NeuroSim’s existing power, performance, and area (PPA) analysis framework, creating a unique environment for systematic CIM accelerator design optimization. The framework achieves runtime only two to three times that of software-only standard GPU inference.
The remainder of this paper is organized as follows: Section II provides technical background on ACIM systems and existing simulation approaches. Section III details our improved framework architecture and generic noise modeling methodology. Section IV presents results from various case studies including inference accuracy analysis of several devices and compute circuits using data from SPICE simulations and tapeout measurements [41, 42, 43, 33, 18, 7]. We also provide design space exploration for a RRAM based ACIM chip to demonstrate the framework’s adaptability and improved running speed. Finally, Section VI discusses implications for future larger-scale AI model benchmarking and concludes the paper.
@@ II-A CIM Array Architecture and Operation
ACIM performs matrix computations by utilizing physical properties of memory arrays. The fundamental building block is the crossbar array, where memory elements are arranged at the intersections of wordlines (WLs) and bitlines (BLs).
This crossbar arrangement is uniquely suited for parallel MAC operations because each row’s input signal (voltage) is simultaneously applied to all memory devices along that row. The resulting outputs (currents or charges) are summed along each column, naturally implementing the dot product. Consequently, large-scale MAC operations can be performed in one analog ‘step’, vastly increasing the throughput compared to digital architectures that must shuttle data back and forth between separate memory and logic units. To date, two primary methods of analog computation have emerged.
Resistive arrays constructed with RRAM, PCM, FeFET that encode neural network weights as conductance states (G), performing MAC operations in the current domain:
@@
where $I_{BL_{i}}$ represents the analog matrix-vector multiplication (MVM) output current on the $i^{th}$ bitline, and $j$ indexes across the wordlines.
Capacitive arrays utilize non-volatile capacitors (nvCap), where the weight is represented by the programmable capacitance of the ferroelectric or charge-trap memory cell [44] [45], thus performing MAC operations in the charge domain:
@@
Alternative charge-based computation approaches utilize fixed capacitors in conjunction with SRAM cells, where the stored digital weight modulates charge accumulation through charge sharing mechanisms [46, 47].
DCIM has emerged as another promising approach [3], where SRAM arrays are augmented with digital multipliers and adder trees to perform integer MAC operations directly within memory. These systems are attractive due to their compatibility with advanced CMOS technology scaling and robust digital operation. However, the digital processing overhead - particularly from the complex adder trees required for partial sum accumulation - impacts array density. Additionally, SRAM leakage power becomes more significant in advanced technology nodes. Both the array scalability and leakage power create challenges for running complete neural networks on-chip.
Each approach is finding its niche, while ACIM remains near the theoretical limit of efficiency for neural network applications. In ACIM designs, the choice of memory technology influences key array characteristics like precision, reliability, and energy efficiency - topics we will explore in detail when discussing noise modeling methodology.
@@ II-B Device and Circuit Level Considerations
The precision and reliability of analog computing operations depend heavily on the properties of individual memory cells. Key device parameters include:
On/Off Ratio: The ratio between maximum and minimum conductance/capacitance states defines the usable dynamic range for weight storage. Finite on/off ratios introduce non-zero current/charge contribution from cells programmed to the off states, which accumulates across array columns and can degrade computation accuracy. This effect becomes particularly significant in larger arrays where hundreds of devices in the off-state contribute to the output. To address this, array designs typically include additional ‘dummy’ columns programmed to off states and are supported in NeuroSim [40], whose outputs are subtracted from the computational columns to cancel the accumulated off-state current/charge. While this technique effectively eliminates the impact of finite on/off ratio, it slightly reduces the available dynamic range for computation since part of the range must be reserved for cancellation. This results in smaller sensing margins and potentially smaller signal-to-noise ratio (SNR). Therefore, high on/off ratios are still a desired characteristic for ACIM memory devices.
Number of States: Supporting high-precision weights requires either multi-level memory cells (MLC) or bit slicing with multiple lower-precision/binary cells. For MLC devices, the number of reliably programmable states determines the weight precision. RRAM, PCM, and nvCap typically achieve 1-4 bits per cell [8, 28, 11], while FeFET have demonstrated >5 bits [14, 48].
At the circuit level, array size is limited mainly by sensing margins, where reliable output detection requires sufficient separation between computational states. This becomes challenging with larger arrays due to increased noise accumulation and reduced signal levels from parasitic effects.
Critical peripheral circuits include:
  • WL/BL drivers to supply the input and output signals.
  • ADCs for digitizing array outputs.
  • Shift-and-add circuits for managing bit-precise computations.
  • Accumulation units on the PE level for partial sum aggregation.
  • Multiplexers for data routing and array access.
The digital circuit blocks in the periphery can be verified through conventional VLSI design methodologies. Therefore, behavioral modeling efforts focus primarily on the analog computational elements.
Figure 2: Procedure for ACIM behavioral simulation in NeuroSim V1.5. (1) Inputs and weights are quantized from floating point to integer. (2) Inputs and weights unfolded and (3) mapped to compute-in-memory arrays. (4) Analog MACs are calculated with analytical circuit models. (5) Noise is injected either to memory devices or to digital MAC outputs. (6) ADC outputs are shifted and added. (7) Outputs from memory arrays are accumulated and concatenated. (8) Accumulated outputs are folded into output tensors. (9) Outputs are de-quantized from integer back to floating-point.
@@ II-C Precision and Data Representation
Neural network parameters must be carefully mapped to CIM arrays while managing precision and computational accuracy. This mapping presents two key challenges: representing multi-bit input activations and storing high-precision weights. For input activations, two main approaches exist:
@@
    Using digital-to-analog converters (DACs) to represent each input value as an analog voltage. While straightforward, this requires complex multi-bit DACs for every row of the array, introducing significant area and energy overhead. The I-V or C-V nonlinearity of the underlying memory devices may also introduces the errors in the input representation.
@@
    Processing each bit of the digital input serially, with outputs shifted and added according to bit significance. This eliminates DAC overhead but requires multiple computational cycles.
@@
    Bit slicing across multiple lower-precision cells. For example, an 8-bit weight can be distributed across cells of various precisions - from a single 8-bit cell to eight 1-bit cells, with intermediate combinations possible.
When slicing the weight among multiple cells, each cell is placed in an adjacent column, with the outputs of each column scaled by its significance post-ADC and added with the other scaled outputs. This operation is referred to as ‘shift-add’ and is described in Fig. 2 and in section III. Using analog shift-add before the ADCs has also been explored [44, 49, 50], but is not explicitly supported in NeuroSim V1.5.
While NeuroSim’s behavioral simulator supports both input representation methods, the PPA estimator only supports the bit-serial approach due to its practical hardware implementation.
In CIM, the fundamental MAC operation is performed in the integer domain, necessitating quantization of the inputs and weights of each layer in the network from floating-point to integer. The analog multiplication and accumulation processes generate integer results, which are digitized through ADCs. Finally, the integer outputs are de-quantized back to floating-point for further processing. Some accelerators explore quantizing the entire network to integers[51, 52], requiring specialized approximations for operations like normalization and activation functions. This approach is particularly applicable to transformers, which require more complex operators like GELU and layer normalization.
@@ II-D Hierarchical System Architecture
The organization of ACIM accelerators stems from fundamental constraints in both device technology and neural network structure. Memory array sizes are typically limited to dimensions between 32×32 and 256×256 cells due to sensing margins and manufacturing considerations. However, neural network layers often require much larger matrix operations. This dimensional mismatch necessitates a hierarchical decomposition of network operations.
The organization of the chip hierarchy in NeuroSim is as follows: At the base level, synaptic arrays perform MAC operations through crossbar structures. These arrays are grouped into processing elements (PEs), which implement weight-stationary, spatial mapping of convolutional operations. For example, when mapping 3D convolutional kernels to 2D array structures, multiple arrays within a PE process different kernel depths in parallel, with their outputs appropriately accumulated.
PEs are further aggregated into computational tiles, each capable of processing a portion of a neural network layer. The tile architecture incorporates buffers for activation storage and accumulation circuits for combining partial results from multiple PEs. This organization enables efficient processing of large layers by distributing computations across multiple tiles while maintaining data locality. At the chip level, multiple tiles operate together through either an H-tree or X-Y bus interconnect structure. This highest level of hierarchy includes global buffers for inter-tile data movement and accumulation circuits for combining results from distributed computations. The entire system implements layer-wise pipelining, where different tiles process consecutive layers of the network simultaneously to maximize throughput.
To balance processing speed between layers of different sizes, weight duplication can be employed within tiles [53]. For instance, if a layer’s weight matrix is smaller than the tile’s capacity, the weights can be duplicated to process multiple input vectors simultaneously, improving throughput.
This hierarchical structure enables flexible mapping of diverse neural network architectures while managing the physical constraints of ACIM arrays. To model this hierarchy, NeuroSim employs analytical models for circuit blocks derived from standard cell libraries in advanced logic nodes. The previously released V1.4 [54] expanded technology support to the leading-edge 1 nm node, introducing stacked nanosheet transistors and studying the scaling path toward a future A5 node with complementary-FET (CFET) technology. Thanks to NeuroSim’s software/hardware co-design philosophy and open-source development model, the tool enables broader research on emerging device technologies, array architectures, on-chip training [55], and advanced packaging [56, 57, 58]. For a detailed analysis of the hierarchy in NeuroSim, readers are referred to [40].

### v2

@@ Abstract
[…]
publicationid: pubid: 0000–0000/00\$00.00 © 2026 IEEE
@@ I Introduction
The rapid growth in AI model size has exposed the importance of energy-efficient processors. Memory access dominates power consumption in AI processing, consuming more than five times the energy of computation [1]. For large models, weights must be loaded periodically from DRAM, shifting the bottleneck from compute capability to memory bandwidth [2].
I-A Compute-in-Memory
Compute-in-memory (CIM) addresses this bottleneck by co-locating multiply-accumulate (MAC) operations with data storage. Digital CIM (DCIM) augments SRAM arrays with multipliers and adder trees [3]. Analog CIM (ACIM) leverages memory cell physics for current-, charge-, or time-domain computation, where weights encode as conductance or capacitance, and column currents/charges are summed to yield MAC results [4, 5]. Various non-volatile memory (NVM) devices enable ACIM arrays, including RRAM [6, 7, 8, 9], PCM [10, 11], FeFET [12, 13, 14], Flash [15], and nvCap [16, 17, 18]. SRAM-based implementations also exist for current-based [19, 20, 21] or charge-based [22, 23] computation.
Both approaches present distinct trade-offs. ACIM with NVM can achieve over 100 TOPS/W [3, 9] and high-density, enabling entire models to be stored on-chip. However, NVM devices suffer from limited write endurance, slow programming, and analog noise from device variations. DCIM offers robust digital computation, unlimited write endurance, and compatibility with advanced CMOS nodes but incurs higher area overhead and lower throughput. These complementary characteristics motivate heterogeneous architectures: ACIM excels for weight-stationary layers where weights are programmed once, while DCIM suits write-intensive operations like attention in transformers. As transformers increasingly dominate AI workloads—from vision transformers (ViTs) to large language models—extending CIM frameworks to support these architectures becomes critical.
I-B CIM Design Challenges
Optimizing CIM accelerators requires navigating a multi-dimensional design space spanning device characteristics (memory technology, precision, on/off ratio), circuit parameters (array dimensions, ADC precision), and system-level choices (quantization schemes, data mapping). The central challenge of ACIM is managing inherently noisy analog computations. Device-level fabrication imperfections cause device-to-device (D2D) variations and stuck-at-faults [24, 25]. Additionally, memory states drift over time [26], and circuit-level thermal noise and process variations degrade ADC sensing margins [27, 18]. At the system level, quantized partial sums compound errors through sequential layers [28, 29]. These multi-layered effects interact in complex ways, necessitating simulation frameworks that can rapidly evaluate both accuracy and hardware metrics across the design space.
I-C Existing Tool Limitations
Direct SPICE simulation is prohibitively slow for large networks, making broad parameter exploration impractical. Existing tools present different trade-offs: CrossSim [30] provides high-fidelity physical simulations but with high computational cost; AIHWKit [31] offers noise modeling primarily for training; CIMLoop [32] focuses on system architectures without accuracy modeling. No existing framework provides integrated accuracy-PPA co-optimization, particularly for trasnformers where self-attention mechanisms introduce unique challenges for CIM deployment.
In this paper, we present NeuroSim V1.5, an integrated simulation framework for systematic co-optimization of accuracy and hardware efficiency in CIM accelerator design. Unlike tools that address device physics or system performance in isolation, NeuroSim V1.5 combines behavioral simulation with detailed hardware analysis to enable comprehensive design space exploration. Our key contributions are:
@@ I-C Existing Tool Limitations
    Support for transformers with hybrid ACIM/DCIM architectures: Our hybrid design augments NeuroSim’s floorplan generator with DCIM tiles used for attention and LUTs for complex activation functions. We conduct a case study on ViT performance using this architecture and compare with CNNs.
@@
    An integrated co-optimization framework combining TensorRT-based quantization, flexible noise modeling validated against SPICE/silicon data, and circuit-level PPA estimation—allowing rapid evaluation of accuracy-efficiency trade-offs across the full design space.
@@
    Expanded device support, including emerging non-volatile capacitive memories (nvCap) and charge-domain compute methods.
@@
[…]
The remainder of this paper is organized as follows: Section II provides technical background on CIM systems. Section III details our framework architecture and noise modeling methodology. Section IV presents case studies including design space exploration, noise sensitivity analysis validated against SPICE and silicon measurements [33, 34, 27, 7], PPA breakdown of Swin Transformer on ImageNet, and detailed error analysis comparing CNN and ViT architectures. Section V discusses future directions.
@@ II-A CIM Array Architecture and Operation
ACIM performs matrix computations by utilizing electrical properties of memory arrays. The fundamental building block is the crossbar array, where memory elements are placed at wordline (WL) and bitline (BL) intersections.
Crossbar arrays enable parallel MACs: input voltages applied to rows produce column currents/charges that naturally sum to implement dot products. This allows large-scale MAC operations in one analog step, vastly increasing throughput. Two primary analog computation methods have emerged.
Resistor-based arrays contain RRAM, PCM, Flash, FeFET or other resistive memory, which encode neural network weights as conductance states (G) and perform MAC operations in the current-domain:
@@
where $I_{BL_{i}}$ is the analog matrix-vector multiplication (MVM) output current on the $i^{th}$ bitline, and $j$ indexes across wordlines.
Capacitor-based arrays contain non-volatile capacitors (nvCap), which encode weights as programmable capacitance [35, 36], performing MAC operations in the charge domain:
@@
Alternative charge-based computation approaches use fixed capacitors with SRAM cells, where stored digital weight modulates charge accumulation via charge sharing [37, 38].
DCIM [3] uses digital multipliers and adder trees for in-memory integer MACs, offering robustness at the cost of higher area overhead compared to ACIM at the same technology node. The main advantage of DCIM is that it can be integrated in any CMOS technology, whereas non-volatile memories are mostly limited to legacy nodes.
@@ II-B Device and Circuit Level Considerations
The precision and reliability of analog computing operations depend heavily on individual memory cell properties. Key device parameters include:
On/Off Ratio: The ratio between maximum and minimum conductance/capacitance states defines the usable dynamic range for weight storage. Finite on/off ratios introduce non-zero current/charge from off-state cells that accumulates across array columns and can degrade accuracy, especially in larger arrays. To address this, designs often include ‘dummy’ columns (supported in NeuroSim [39]) whose outputs are subtracted to cancel accumulated off-state current/charge. While effective, this slightly reduces computation dynamic range, resulting in smaller sensing margins and potentially smaller signal-to-noise ratio (SNR). Thus, high on/off ratios remain desirable for ACIM devices.
Number of States: Supporting high-precision weights requires either multi-level memory cells (MLC) or bit slicing with multiple lower-precision/binary cells. For MLC devices, the number of reliable programmable states determines weight precision. RRAM, PCM, Flash, and nvCap typically achieve 1-4 bits per cell [8, 40, 11], while FeFETs have demonstrated >>5 bits [13, 41].
At the circuit level, array size is limited mainly by sensing margins, as reliable output detection requires sufficient separation between analog states. This becomes challenging with larger arrays due to increased noise accumulation and reduced signal levels from parasitics. Peripheral circuits include WL/BL drivers, ADCs, shift-and-add circuits, inter-array accumulation units, and multiplexers for data routing. Peripheral digital circuit blocks are verifiable via conventional VLSI methods; thus our behavioral modeling focuses primarily on analog computational elements.
@@ II-C Precision and Data Representation
Neural network parameters must be carefully mapped to CIM arrays, managing precision and computational accuracy. This presents two key challenges: representing multi-bit input activations and storing high-precision weights. For input activations, three main approaches exist:
@@
    Multi-bit DACs convert inputs to analog voltages but require significant area/energy and may introduce device non-linearity errors.
@@
    Bit-serial processing eliminates DAC overhead at the cost of multiple computational cycles.
  3. 3.
    Time-encoding [42, 43, 44] offers simplified ADCs at the cost of latency. Time-domain operation is not currently supported in NeuroSim.
@@
    Bit Slicing: Used when desired weight precision ($b_{w}$) exceeds memory cell bit precision ($b_{cell}$). E.g., an 8-bit weight can be split across multiple memory cells.
  2. 2.
    Direct Mapping: Used when weight precision ($b_{w}$) is less than or equal to cell precision ($b_{cell}$). In this case, $N_{cell}=1$, and the entire analog weight maps directly to a single memory cell’s state.
Bit-sliced weights occupy adjacent columns; outputs are scaled and summed post-ADC (’shift-add’, Fig. 2). The PPA estimator models bit-serial inputs for hardware accuracy. Pre-ADC analog shift-add for multi-bit weights [45, 46] and pre-ADC input accumulation schemes [47, 48] can reduce compute cycles due to bit-slicing. While we have explored these techniques [35], they are not explicitly supported in this version but can be added with minimal user modification.
CIM MACs operate in integer domain, requiring float-to-int quantization. ADCs digitize analog results, which are then de-quantized to floating-point. Some accelerators explore quantizing the entire network to integers [49, 50], requiring specialized approximations for operations like normalization and activation functions.
[…]
@@ II-D Hierarchical System Architecture
CIM accelerator organization stems from device technology and neural network structure constraints. Memory array sizes are typically $32\times 32$ to $256\times 256$ cells due to sensing margins and manufacturing yields. However, neural network layers often require much larger matrix operations, necessitating hierarchical decomposition.
NeuroSim’s chip hierarchy: At base level, crossbar arrays compute partial sums of larger linear/convolutional layers. Crossbars are grouped into processing elements (PEs), where partial sums are accumulated. PEs are then grouped into tiles which include activation buffers and accumulation circuits for combining PE results– enabling efficient large layer processing by distributing computations while maintaining data locality. At chip level, multiple tiles operate via an H-tree or X-Y bus interconnect. This highest level includes global buffers for inter-tile data movement and accumulation circuits for distributed computation results. The system implements layer-wise pipelining, where different tiles process consecutive network layers simultaneously to maximize throughput.
NeuroSim assumes all weights fit on-chip in NVM, enabling a fully weight-stationary dataflow. This design avoids frequent reprogramming that would stress NVM write endurance and increase latency. Small models that would underutilize a design’s compute capacity are duplicated to boost throughput [51]. For models exceeding single-die capacity, we have developed separate frameworks for 3D chiplet integration [52].
The PPA estimator generates hardware configurations optimized to a specific network architecture (e.g. CNNs). While the floorplan is generated based on one selected network architecture, the hardware can execute multiple models of the same type through runtime reconfigurable weight mapping and dataflow control, as demonstrated in our prior work [53].
To simulate latency, area, and energy consumption, NeuroSim uses analytical models (C++) of circuit blocks based on standard cell libraries in advanced logic nodes. V1.4 [54] expanded technology support to 1nm, including stacked nanosheet transistors, as well as support for future A5 node with complementary-FETs. NeuroSim’s software/hardware co-design philosophy and open-source model enable research on emerging devices, array architectures, on-chip training [55], and advanced packaging [52, 56, 57]. For a more detailed hierarchy analysis, see [39].
Figure 2: Procedure for CIM behavioral simulation in NeuroSim V1.5. (1) Inputs and weights are quantized from floating point to integer. (2) Inputs and weights unfolded and (3) mapped to compute-in-memory arrays. (4) MACs are calculated with analytical circuit models. (5) Noise is injected either to memory devices (pre-ADC) or to digital MAC outputs (post-ADC). (6) ADC/Adder tree outputs are shifted and added. (7) Outputs from memory arrays are accumulated and concatenated. (8) Accumulated outputs are folded into output tensors. (9) Outputs are de-quantized from integer back to floating-point.